Identify and trace the conicoid y^2+z^2 = x. Describe its sections by the planes
x = 0;y = 0 and z = 0.
1
Expert's answer
2019-03-18T10:31:36-0400

It is an elliptic paraboloid:



sections by the planes

x = 0: it is only one point x=0.y=0,z=0

sections by the planes

y = 0: z^2 = x - it is parabola in x - z plane with x axis of symmetry

sections by the planes

z = 0: y^2 = x - it is parabola in x - y plane with x axis of symmetry